# Simulating Vibrational Dynamics on Bosonic Quantum Devices

### v1

@@ II Theory
Here, we use atomic units with $\hbar=1$, and have truncated the Taylor expansion of the vibrational potential energy surface $V(\bm{q})$ to fourth order around the equilibrium geometry. $V^{\{3\}}_{ijk}$ and $V^{\{4\}}_{ijkl}$ are cubic and quartic coefficients respectively. The position and momentum operators follow canonical commutation relations $[q_{j},q_{k}]=[p_{j},p_{k}]=0$, and $[q_{j},p_{k}]=i\delta_{jk}$. The vibrational Hamiltonian can also be expressed as a function of bosonic creation and annihilation operators,
@@ II.3 Constructing the Propagator from Hamiltonian Fragments
The unitary operators corresponding to the diagonal fragments can also be decomposed into gates available on bosonic hardware. The unitary $e^{i\tilde{H}_{0}t}$ corresponding to the quadratic diagonal fragment $\tilde{H}_{0}$ can be written as a product of phase space rotation gates and a phase factor,
@@ III.1 Tunneling Dynamics in Tropolone
Here $x$ and $y$ are dimensionless coordinates representing the high-frequency proton tunneling mode and the nearly-planar hydrogen bond stretching mode respectively wojcik2009 . Figure 1 depicts a contour plot of the potential. The parameters in the potential are $\omega_{x}=3594$ cm-1, $\omega_{y}=414$ cm-1, $x_{0}=3.15$ and $\alpha=0.301$. Solvable fragments for this double well Hamiltonian are found using the GFRO algorithm presented in Sec. II.2, with Powell’s algorithm Powell1964 used for the non-linear optimization. A tolerance of $10^{-2}$ cm-1 is used in step 4 of the GFRO algorithm to stop fragmentation, resulting in a total of four solvable fragments including one quadratic fragment.
Figure 3: The probability density $\rho(x,y,t)$ is plotted as a function of time for the propagation of the initial Gaussian wavepacket |$\ket{\psi(0)}$ defined in Eq. (17). The propagator is constructed using solvable fragments for the double well Hamiltonian using a Trotter time step of 0.01 a.u. The maximal evolution time $\tau$ is defined in the text. Gray lines denote contours of the potential to indicate locations of the wells.
We track tunneling dynamics of a Gaussian wavepacket initially centered at the bottom of the left well (first panel in Figure 3). It is created by displacing the vacuum state |$\ket{\mathbf{0}}$ to be centered at the potential minimum located at $x=-x_{0},y=0$,
@@
The average energy of this wavepacket is 4429 cm-1, well below the barrier height of 14042 cm-1. Owing to its significant overlap with the ground and first excited eigenstates of the potential, the wavepacket is expected to coherently tunnel through the barrier on a timescale close to the period corresponding to the tunneling splitting between these two states ($\tau$). To construct the propagator, we find Hamiltonian fragments and use the Trotter approximation. Figure 2 presents the overlap error as a function of the Trotter step, with
@@
where |$\ket{\psi_{ex}(\tau)}$ and |$\ket{\psi_{tr}(\tau)}$ are the wavefunctions propagated exactly and using the Trotter approximation respectively. A time step of $0.01$ a.u. is chosen to keep the overlap error less than $10^{-2}$. For larger systems, well established methods to estimate Trotter approximation errors can be used to obtain the appropriate step size childs2021 ; martinez2023 ; rendon2024 ; martinez2024 . Figure 3 plots the probability density $\rho(x,y,t)$ as a function of time, propagated using the Hamiltonian fragments found for the double well Hamiltonian. As expected, the wavepacket coherently tunnels through the barrier by $t=\tau/2$ and then tunnels back by $t=\tau$. Noting that the propagator is constructed from solvable fragments that can be implemented on a bosonic device, this demonstration offers a methodology to simulate tunneling dynamics on currently available devices.
@@ III.2 Vibrational Eigenenergies
To demonstrate our approach, we calculate the eigenenergies of $H_{\rm{eff}}$ for the small molecules considered here. The maximal evolution time is generally decided on by the accuracy required in the eigenvalue calculation. Furthermore, the total evolution will need to be partitioned into smaller time steps to ensure that the Trotter approximation error is maintained below the acceptable error threshold. Here, for the purpose of demonstrating our fragmentation scheme, we use an evolution time of $t=1$ a.u., which is sufficiently small to obtain Trotter approximation errors in eigenvalues $<1$ cm-1. Other methods based on perturbative estimates for the Trotter approximation error for eigenvalues may also be employed martinez2024 . Table 2 compares the first four eigenenergies for the molecules considered here obtained using the Trotter approximation to those obtained by the exact diagonalization of the Hamiltonian. For all molecules considered, the fragmentation procedure yields errors $<1$ cm-1 compared to the exact results.
@@ IV Conclusion
The fragmentation scheme is tested on two applications. We perform coherent tunneling dynamics for a two-dimensional double-well potential. Noting that this model Hamiltonian can be partitioned into just four solvable fragments, this system may serve as a promising candidate for an experimental demonstration of our approach in the future studies. Furthermore, the ability of the approach to fragment Hamiltonians with higher number of degrees of freedom is demonstrated by calculating vibrational eigenenergies for molecules with up to four vibrational modes. A comparison of the number of bosonic fragments with the number of fully commuting Pauli fragments highlights that the bosonic fragmentation scheme is $13-17$ times cheaper in terms of fragment counts. Furthermore, we wish to emphasize that for systems where a more accurate description of the potential energy surface is required, our fragmentation scheme can also be modified to the vibrational Hamiltonian written in the second quantized Christiansen form christiansen2004 that is based on the $n-$mode representation of the potential bowman2003 . Our approach extends the range of applicability of bosonic quantum devices to simulations of chemical dynamics and spectroscopy.
@@ Appendix C Bloch-Messiah Decomposition of the Bogoliubov Unitary
Combining all these elements, the Bogoluibov unitary can be written as $\mathcal{U}_{b}=\mathcal{U}_{d}\,\mathcal{U}_{q}=\mathcal{U}_{d}\,\mathcal{U}_{3}\,\mathcal{U}_{2}\,\mathcal{U}_{1}$. This is highlighted in Eq. (10).

### v2

@@ II Theory
Here, we use atomic units with $\hbar=1$, and have truncated the Taylor expansion of the vibrational potential energy surface $V(\bm{q})$ to fourth order around the equilibrium geometry. $N$ is number of vibrational degrees of freedom, and $V^{\{3\}}_{ijk}$ and $V^{\{4\}}_{ijkl}$ are cubic and quartic coefficients respectively. The position and momentum operators follow canonical commutation relations $[q_{j},q_{k}]=[p_{j},p_{k}]=0$, and $[q_{j},p_{k}]=i\delta_{jk}$. The vibrational Hamiltonian can also be expressed as a function of bosonic creation and annihilation operators,
@@ II.3 Constructing the Propagator from Hamiltonian Fragments
The unitary operators corresponding to the diagonal fragments can also be decomposed into gates available on bosonic hardware. The unitary $e^{iD_{0}t}$ corresponding to the quadratic diagonal fragment $D_{0}$ can be written as a product of phase space rotation gates and a phase factor,
@@ III.1 Tunneling Dynamics in Tropolone
Here $x$ and $y$ are dimensionless coordinates representing the high-frequency proton tunneling mode and the nearly-planar hydrogen bond stretching mode respectively wojcik2009 . Figure 1 depicts a contour plot of the potential. The parameters in the potential are $\omega_{x}=3594$ cm-1, $\omega_{y}=414$ cm-1, $x_{0}=3.15$ and $\alpha=0.301$. Solvable fragments for this double well Hamiltonian are found using the GFRO algorithm presented in Sec. II.2, with Powell’s algorithm Powell1964 used for the non-linear optimization. A tolerance of $10^{-2}$ cm-1 is used in step 4 of the GFRO algorithm to stop fragmentation, resulting in a total of four solvable fragments including one quadratic fragment. We track tunneling dynamics of a Gaussian wavepacket initially centered at the bottom of the left well. It is created by displacing the vacuum state |$\ket{\mathbf{0}}$ to be centered at the potential minimum located at $x=-x_{0},y=0$,
@@
The average energy of this wavepacket is 4429 cm-1, well below the barrier height of 14042 cm-1. Owing to its significant overlap with the ground and first excited eigenstates of the potential, upon evolution, the wavepacket coherently tunnels through the barrier on a timescale close to the period corresponding to the tunneling splitting between these two states ($\tau$). To construct the propagator, we find Hamiltonian fragments and use the Trotter approximation. Figure 2 presents the overlap error as a function of the Trotter time step, with
@@
where |$\ket{\psi_{ex}(\tau)}$ and |$\ket{\psi_{tr}(\tau)}$ are the wavefunctions propagated exactly and using the Trotter approximation respectively. The trotter time step can be chosen to keep the overlap error below a chosen threshold. For larger systems, well established methods to estimate Trotter approximation errors can be used to obtain the appropriate step size childs2021 ; martinez2023 ; rendon2024 ; martinez2024 . Noting that the propagator is constructed from solvable fragments that can be implemented on a bosonic device, this demonstration offers a methodology to simulate tunneling dynamics on currently available devices.
@@ III.2 Vibrational Eigenenergies
To demonstrate our approach, we calculate the eigenenergies of $H_{\rm{eff}}$ for the small molecules considered here. The maximal evolution time is generally decided by the accuracy required in the eigenvalue calculation. Furthermore, the total evolution will need to be partitioned into smaller time steps to ensure that the Trotter approximation error is maintained below the acceptable error threshold. Here, for the purpose of demonstrating our fragmentation scheme, we use an evolution time of $t=1$ a.u., which is sufficiently small to obtain Trotter approximation errors in eigenvalues $<1$ cm-1. Other methods based on perturbative estimates for the Trotter approximation error for eigenvalues may also be employed martinez2024 . Table 2 compares the first four eigenenergies for the molecules considered here obtained using the Trotter approximation to those obtained by the exact diagonalization of the Hamiltonian. For all molecules considered, the fragmentation procedure yields errors $<1$ cm-1 compared to the exact results.
@@ IV Conclusion
The fragmentation scheme is tested on two applications. We perform coherent tunneling dynamics for a two-dimensional double-well potential. Noting that this model Hamiltonian can be partitioned into just four solvable fragments, this system may serve as a promising candidate for an experimental demonstration of our approach in future studies. Furthermore, the ability of the approach to fragment Hamiltonians with higher number of degrees of freedom is demonstrated by calculating vibrational eigenenergies for molecules with up to four vibrational modes. A comparison of the number of bosonic fragments with the number of fully commuting Pauli fragments highlights that the bosonic fragmentation scheme is $13-17$ times cheaper in terms of fragment counts. Furthermore, we wish to emphasize that for systems where a more accurate description of the potential energy surface is required, our fragmentation scheme can also be modified to the vibrational Hamiltonian written in the second quantized Christiansen form christiansen2004 that is based on the $n-$mode representation of the potential bowman2003 . Our approach extends the range of applicability of bosonic quantum devices to simulations of chemical dynamics and spectroscopy.
@@ Appendix C Bloch-Messiah Decomposition of the Bogoliubov Unitary
Combining all these elements, the Bogoliubov unitary can be written as $\mathcal{U}_{b}=\mathcal{U}_{d}\,\mathcal{U}_{q}=\mathcal{U}_{d}\,\mathcal{U}_{3}\,\mathcal{U}_{2}\,\mathcal{U}_{1}$. This is highlighted in Eq. (10).